A ship keel plate is 5.25m x 1.85m. Determine the maximum force on the plate when the ship is in fresh water at draught (height) of 6.14m.



1
Expert's answer
2021-01-26T08:41:21-0500

"F=p\cdot A=\rho gh\cdot A=1000\cdot9.8\cdot 6.14\cdot5.25\cdot1.85="


"= 584421(N)\approx584(kN)" . Answer